OpenStudy (anonymous):

Explain in your own words how to evaluate a polynomial for a given value of the variable. Demonstrate the process with an example.

OpenStudy (chaise):

Simply substitute in your value your are supplied with. For example: x^3 + 2x + 18 = ? For values of x = 3 x^3 + 2x +18 = 3^3 + 2x3 + 18 = 51 I hope this helped and was what you were after.

OpenStudy (chaise):

For some reason my enter button doesn't work so it is must up. Erm... I hope you understood this.

OpenStudy (anonymous):

yes chaise thank you.